How do you solve for x when the problem is: x/2 - 1/3 = 3/2?
Found 3 solutions by lynnlo, ikleyn, josgarithmetic:
Answer by lynnlo(4176)
Answer by ikleyn(53304)

.
How do you solve for x when the problem is: x/2 - 1/3 = 3/2?
~~~~~~~~~~~~~~~~~~~~~~~~~~

Multiply all the terms by 2.  You will get

    x - 2/3 = 3.


Now add 2/3 to both sides to get

    x = 3 2/3.


ANSWER.  x = 3 2/3,  or  x = 11/3.

Solved.

The answer in the post by @lynnlo is incorrect.
For the peace in your mind, ignore his post.



Answer by josgarithmetic(39670)